A student placed 20.0 g of glucose (C6H12O6) in a volumetric flask, added enough water to dissolve the glucose by swirling, then carefully added additional water until the 100. mL mark on the neck of the flask was reached. The flask was then shaken until the solution was uniform. A 40.0mL sample of this glucose solution was diluted to 0.500 L. How many grams of glucose are in 100. mL of the final solution?
Answer #1

The initial solution concentration is calculated:

C1 = g / mL = 20 g / 100 mL = 0.2 g / mL

Through the dilution ratio, the final concentration is calculated:

C2 = C1 * V1 / V2 = 0.2 g / mL * 40 mL / 500 mL = 0.016 g / mL

The glucose mass is calculated in 100 mL of solution:

g = 0.016 g / mL * 100 mL = 1.6 g

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.